OpenStudy (eric_d):

Differentiate following with respect to x 23) 1/x ln x

OpenStudy (eric_d):

@ganeshie8

OpenStudy (freckles):

\[\frac{1}{x}\ln(x) \text{ or } \frac{1}{x \ln(x) }\]

OpenStudy (eric_d):



OpenStudy (jhannybean):

\[d(\left(\frac{1}{x} \ln(x)\right)\]\[f=\frac{1}{x} \implies x^{-1}\]\[g=\ln(x)\]

OpenStudy (surry99):

product rule

OpenStudy (jhannybean):

So product rule states: \(f'g + g'f\)\[d(f(x))=(x^{-1})'(\ln(x)) + (x^{-1})(\ln(x))'\]

OpenStudy (jhannybean):

Can you take it from here

OpenStudy (eric_d):

Trying to understand...

OpenStudy (anonymous):

Hmmm.. This is easy!

OpenStudy (jhannybean):

Take the derivative of what i labeled as \(f\) and multiply it by function \(g\) , then take the derivative of \(g\) and multiply it by \(f\)

OpenStudy (anonymous):